OpenStudy (anonymous):

What would the function rule be if x= 10,11,12,13 and y= 200, 202, 204,206

OpenStudy (anonymous):

(10,200) (11,202) (12,204) (13,206)

OpenStudy (anonymous):

oh sorry i just read over your question and i missed one word, rule.

OpenStudy (anonymous):

its's alright

OpenStudy (anonymous):

Okay so if you have the same number twice in the x column it's not a function

OpenStudy (anonymous):

does that help any?

OpenStudy (anonymous):

yes, thank you

OpenStudy (anonymous):

Welcome, glad to help.

OpenStudy (amoodarya):

look those number they grow linear it means they are on the line but for calculate suppose f(x)=ax+b now put (10,200) and (11,202) and find a,b

OpenStudy (amoodarya):

y=ax+b (10,200) (11,202) 200=10a+b 202=11a+b solve the system of equation

OpenStudy (amoodarya):

f(x)=2x+180